OpenStudy (anonymous):

Maritta was given a raise of $0.75 an hour, which brought her hourly wage to $12.25. write and solve an equation to determine maritta's hourly wage before her raise

OpenStudy (mitu12):

take the new hourly wage and subtract the raise from it to get what the old hourly wage was..to see how much it increased... 12.25 - 0.75 = 11.5 $11.50 was the old hourly wage.

OpenStudy (anonymous):

If Marietta was given a raise of $0.75 an hour, which brought her hourly wage to $12.25, her hourly wage was $11.50 ($12.25 - $0.75) before the increase. $11.50 + $0.75 = $12.25. $12.25 is 6.52174% larger than $11.50. So it was a 6.5% raise. An equation is a statement that the values of two mathematical expressions are equal: 11.5 + 0.75 = 12.25. 11.5+0.75 = 12.25 - 0.75.

OpenStudy (anonymous):

got it thanks for help @Informants and @mitu12

OpenStudy (anonymous):

No problem heres more work if yu need it :) 12.25 - 0.75 = x x = 11.50 <--- That is the x <3

OpenStudy (mitu12):

no problem